# Magnetic field amplification and decay in cosmic string wakes

### v2

@@ 2.2 Magnetic fields in rotational flows
We now consider the role of the magnetic field in the simulation. The magnetic field $B$ has two components (in $x$-and $y$-direction). For a uniform magnetic field, we give fixed values to both components. Generally, the generation of vorticity in magnetohydrodynamics leads to turbulence in the plasma. However, for a turbulence to develop, we need to introduce a constant source for generating vortices at different length scales. As mentioned before, we have a constant background magnetic field in the simulation so the vorticity is generated at a fixed lengthscale. For turbulence, the plasma has to be non-ideal and the net vorticity need not be conserved. In our simulation, we do not introduce any other potential so the lengthscales of the generated vortices are fixed by the scale of the cosmic string wake and the background magnetic field. Hence no turbulence occurs in our system. So, the dynamo mechanism is completely absent from our simulations.
@@
We plot the magnetic field lines in the simulation. As expected, the magnetic field lines are deformed due to the flow of the fluid. We calculate the total magnetic field in the wake region as well as the peak value of the field. As long as the magnetic field remains constant, we do not expect any change in the magnetohydrodynamic flow. The flow remains steady with minor changes due to the velocity perturbation. We proceed to give a perturbation to the magnetic field. Since we would like the field to be amplified, we give a sheared perturbation to the magnetic field as it is previously known that sheared magnetic fields may lead to amplification of the magnetic field king . The field is perturbed first along the $x$- direction and then in the $y$-direction, the perturbations are denoted by $B_{x}$ and $B_{y}$, respectively. The results of these perturbations are discussed in the next section.
3 Sheared magnetic fields and field amplification
@@
For a frozen in magnetic field, the frozen in condition ensures that the magnitude of the magnetic field will be related to the density of the plasma. It has been shown that an isotropic collapsing plasma can lead to the amplification of a magnetic field even in the absence of a dynamo mechanism king . The argument was that the flux in the frozen in plasma is conserved, so in the collapsing region, as the cross-sectional area decreases, the magnitude of the magnetic field increases. Now, the formation of a wake behind a cosmic string leads to a substantial increase in the density of plasma behind the cosmic string. In this case too, there is the possibility that the magnitude of the magnetic field will be enhanced.
@@
To understand this, we looked at two different kinds of magnetic field perturbations. We start with an exponentially decreasing magnetic field perturbation. The strength of the magnetic field is proportional to the density of the magnetic field lines. As mentioned before, the cosmic string is moving in the $x$-direction. Due to the velocity perturbation, the fluid velocity has a component in the $y$-direction, and therefore, the magnetic field lines feel a force along the $y$-direction. Therefore, we start with a sheared perturbation in the direction of motion of the cosmic string. Fig.2 represents the graph of the total magnetic field with time for a perturbation of $B_{x}=B_{0}e^{-\alpha_{1}|y|}$. We call this a sheared magnetic field as the $x$ component of the magnetic field ($B_{x}$) at any given point has a variation in the $y$-direction. The lengthscale of the variation is given by the parameter $\alpha_{1}$. We have used $B_{0}=1$.
@@
Fig.3 represents the graph of the peak magnetic field evolution with time for $\alpha_{1}=0.01$. As the graphs shows there is an increase in the peak magnetic field value which will lead to the total magnetic energy being enhanced. The figure shows that the peak magnetic field value becomes at least five times the initial value. The peak magnetic value shown on the $y$- axis is the maximum value of the magnetic field at a specific point on the grid. Similarly, the total energy also becomes five times the initial energy value, the $|B|^{2}$ on the y - axis is the summation of the magnetic energy density over all the grid points. Together, they indicate that the magnetic field value not only increases at a particular point, it increases over the whole of the cosmic string wake. So, we see that the net magnetic field gets amplified without any external dynamo mechanism. We now discuss in detail why this is happening.
There is a velocity perturbation on the particles along the $y$-direction, so the particles move towards the conical space behind the cosmic string. We have given a sheared magnetic field as the magnetic field perturbation. Though we are perturbing the $x$ component, the variation is along the $y$-direction. As the magnetic field lines are being pushed in the $y$-direction by the velocity flow, so our perturbation is enhancing the effect of the compression of the magnetic field lines due to the fluid flow. We find that we get an amplification of the field for all values of $\alpha_{1}$.
The amplification of the magnetic field without any external turbulent forces can be attributed to the deformation tensor of the fluid element. The deformation tensor of the fluid element has been used to explain the amplification of frozen in magnetic fields in ref king . The deformation parameter $D_{ij}$ represents the deformation of an infinitesimal fluid element of initial side length $\delta q_{j}$, which in this case is deformed due to the deficit angle in the local space of the cosmic string. The deformation tensor is defined by
@@
Fig.4 is used to explain how the magnetic field gets amplified for this given magnetic field perturbation. As has already been pointed out in ref king , the magnitude of the magnetic field is affected only when the fluid motion is perpendicular to the magnetic field lines. In this case, for any given value of $x$, the magnetic field varies along $y$ as an exponentially decaying field. This is illustrated in Fig.4. The $q_{1}$ in the figure is the $x$-axis, while the $q_{2}$ is the $y$-axis. The angle $\theta$ by which the fluid element is deformed is proportional to the deficit angle of the cosmic string. The comoving volume element is preserved after the deformation, so $\frac{V^{fin}}{V^{ini}}=D_{ij}$. Since the volume can be related to the density of the fluid, we will have
@@
where $A$ denotes the initial and final cross-sectional areas through which the magnetic field line passes before and after deformation. If the length perpendicular to this area is referred to as $L$, from volume conservation we will be have,
@@
The angle between $dq_{2}$ and $L^{fin}$ is the deficit angle $\theta$. This vector $L^{fin}$ has not been shown on the Fig.4 as it is rather difficult to depict in this case. Thus we will have
@@
This will mean that $B\propto\rho^{1/2}$. Similarly in any case, that is considered, the final magnetic field is determined by the final density of the deformed fluid.
Now, in the wake region, we will always have $\rho^{fin}>\rho^{ini}$, therefore the magnetic field will always be amplified in the wake of the cosmic string due to the deformation of the field lines in the presence of the deficit angle. This will happen as long as the magnetic field remains frozen in the plasma. We have checked the results for different values of $\alpha_{1}$ and in all the cases we see an increase in the magnitude of the magnetic field in the wake region.
The magnetic field in the cosmic string wake can have components in both the $x$-and $y$-direction. So, our next case would be a perturbation of $B_{y}=B_{0}e^{-\alpha_{2}|(x-800)|}$. The origin is shifted to get a better figure of the final wake. It does not have anything to do with the field magnitude. Here, we depict the lengthscale by the parameter $\alpha_{2}$. The initial value is still $B_{0}=1$.
@@
Figure 6: Evolution of the peak value of the magnetic field in the cosmic string wake. The value satuarates to about five times the initial value even in this case.
As is seen from Fig.5 and Fig.6, we still get a amplification of the magnetic field. The maximum value is approximately five times the initial value again. In this case too, the field lines are deformed by the perturbation. As we have taken the flow velocities and the perturbations along the $x$-and $y$-axes, the amplification is similar in both the cases. If the axes were not aligned we might have got different results for the amplification. In both our cases the off-diagonal elements which are non-zero are driven by symmetry consideration. So, we can show that the magnetic field is proportional to the density of the wake in both the cases. We have repeated the simulation with a different periodic magnetic field and obtained similar results about the magnetic field amplification. For simplicity, we have taken one of the axes to be aligned with the flow, even if that is not the case, the amplification will still happen, only the symmetry consideration cannot be used to calculate the off-diagonal elements.
4 Magnetic field decay
@@
So, for a perturbation perpendicular to the flow velocity of the plasma, the lengthscale of the perturbation determines whether the field decays or it gets amplified. As discussed in the previous section, the Alfven’s theorem leads to the frozen in magnetic field lines. Due to the deformation of the fluid, the magnetic field lines get squeezed, and the amplification of the field is observed. However, in this case too the fluid is deformed but the magnetic field does not get amplified. This seems to indicate that Alfven’s theorem is no longer valid for this lengthscale.
@@
In the fluid flow of the wake, the magnetic field lines initially move with the velocity of the fluid. The advection velocity of the loops are then given by the velocity of the particles in the plasma. Assuming the majority of the particles to be electrons, one can then calculate the gyroradius of the plasma electrons. The gyroradius is given by $r_{g}=\frac{mcv_{y}}{qB}$ where $v_{y}$ is the perpendicular component of the velocity of the electron. We find that the gyroradius gives us a good estimate of whether $l_{0}$ has been reached. Based on the fluid velocity, we calculate an approximate value of the electron gyroradius $r_{g}$. We find that the magnetic field gets amplified only when the particle gyroradius $r_{g}$ is smaller than the scale at which the magnetic field perturbation varies, i.e. for $r_{g}<1/\alpha_{2}$. Since in our numerical simulations the value of $r_{g}\sim 1.31$, hence we find that when the value of $\alpha_{2}\sim 1$, the field decays instead of getting amplified.
@@
The relation of the lengthscale $l_{0}$ to the gyroradius can be obtained from the trajectory of charged particles in non-uniform magnetic fields. Previously, the motion of charged particles has been studied in specific sheared force-free magnetic fields vekstein . In this case, we have a non-uniform sheared magnetic field perturbation perpendicular to the flow direction of the fluid. The motion of the electrically charged particles in a perturbed magnetic field has been studied using the approximate description called the guiding center or drift approximation northrop . The charged particle motion in the plasma is decomposed into a nearly circular fast gyromotion about a guiding center and the trajectory of the guiding center itself. Neither of the trajectories are constant for the non-uniform magnetic field. The overall motion is obtained by solving the second-order differential equations for the guiding center. Usually, for better understanding, the equations are separated into parallel and perpendicular components of the given magnetic field. In our present case, the perpendicular direction is the direction along which the field is non-uniform. The velocity in the perpendicular direction (i.e. along the non-uniform direction) is then given by a gradient drift ($v_{g}$) and a curvature drift ($v_{c}$). These are defined by,
@@
5 Summary and Conclusions
We have done a detailed study of the amplification of magnetic fields in the wake of a cosmic string in the absence of a dynamo mechanism. Though hydrodynamic wakes and shocks are well studied for a cosmic string, it is only in recent times that magnetohydrodynamic shocks have also been studied in the cosmic string wake. Magnetohydrodynamic shocks are difficult to study as the magnetic field lengthscales play a major role in the generation and evolution of the shock dahlburg . Among cosmic string wakes, MHD shocks have been discussed previously in ref. beresnyak . In recent times, more detailed simulations of such wakes have been carried out in ref. soumen . One of the crucial factors that were never considered in these studies was the lengthscale of the perturbations that affect the magnetic and density fields. Lengthscales of both density inhomogeneities as well as magnetic fields affect structure formation and later evolution of large scale structures in the universe kim . For cosmic strings, this is particularly important as density perturbations depend on the small-scale structure of the strings wu as well as the nature of the particle motion around them abhisek .
@@ 5 Summary and Conclusions
In this respect, our study is important as it clearly shows the effect of perturbations of different lengthscales on the magnetic field. We find that the magnetic field gets amplified by the presence of sheared perturbations close to the core of the cosmic string. We also find that the amplification strongly depends on the nature of the magnetic field and it’s lengthscale. The amplification of the magnetic field occurs due to the squeezing of the magnetic flux in the wake of the cosmic string. This can only happen when the magnetic field lines are frozen into the plasma. The shear causes a deformation in the fluid. Assuming the wake is symmetric about the flow axis, we show that the deformation tensor will have non-zero off-diagonal terms. This leads to the fact that the magnetic field strength will become proportional to the square root of the fluid density. Depending on the orientation of the deformation, the magnetic field can also be proportional to the density. Since the overdensity in the wake is always greater than the initial density of the flow, the magnetic field flux gets amplified in all the different cases.
@@
We then proceed to show that the amplification also depends on the nature of the perturbation of the magnetic field. For a sheared perturbation which is perpendicular to the direction of flow of the fluid, the magnetic field lengthscale determines whether the magnetic field will amplify or decay. This is because below a certain lengthscale the Alfven’s theorem of frozen in magnetic field may not be valid and local interactions will determine the dissipation of the magnetic field. We have determined that the length scale below which the magnetic field decays is determined by the gyroradius of the particles being considered. If the lengthscale of the magnetic field is much larger than the gyroradius, Alfven’s theorem remains valid and we get the amplification of the field. However, if the lengthscale of the magnetic field perturbation is of the same order or less than the gyroradius of the electron then the field decays.
@@
The magnetic field due to the Biermann mechanism is due to the small-scale inhomogeneities generated in the cosmic string wake. Since the density perturbations of a long cosmic string are of varied lengthscales based on how they are generated, it is important to understand the behaviour of the magnetic field for perturbations of different lengthscales. There are cosmic strings with small-scale structures on them which can generate small-scale non-linear density perturbations. It is also known that non-Gaussian density perturbations can be generated by long strings dacunha . As long as Alfven’s theorem holds, the magnetic field configurations will directly relate to the density perturbations. In this work, we have shown that the lengthscale of the magnetic field perturbations will determine if Alfven’s theorem is applicable and we will have a magnification of the magnetic field. We have also seen that for magnetic perturbations with lengthscales of the order of the gyro radius or less, Alfven’s theorem breaks down. This in turn can lead to the possibility of magnetic reconnection in the cosmic string wake. Since, in recent times, magnetic reconnection has been seen as a viable possibility in the narrow wakes of the cosmic string, this study will lead to a further understanding of the evolution of magnetic fields in cosmic string wakes.
@@
Our study has some limitations as it is constrained to two dimensions only. The current study therefore has one dimensional sheared magnetic fields only. The actual wake structure is three-dimensional, so it is also possible that the perturbation is a planar sheared field. The evolution of such perturbations may lead to further interesting results. We have also considered ideal MHD only, it is quite possible that the reconnection possibility and the breakdown of Alfven’s theorem will be different for resistive MHD. In fact, there might be a higher probability of obtaining magnetic reconnection in that case. These and other aspects of magnetic field evolution in cosmic string wakes need to be studied in more details in the future.

### v3

@@ II.2 Magnetic fields in rotational flows
We now consider the role of the magnetic field in the simulation. The magnetic field $B$ has two components (in $x$-and $y$-direction). For a uniform magnetic field, we give fixed values to both components. Charged particles in magnetic field move in rotational motion and generate vorticity. Generally, the generation of vorticity in magnetohydrodynamics leads to turbulence in the plasma. However, for a turbulence to develop, we need to introduce a constant source for generating vortices at different length scales. As mentioned before, we have a constant background magnetic field in the simulation so the vorticity is generated at a fixed lengthscale. For turbulence, the plasma has to be non-ideal and the net vorticity need not be conserved. In our simulation, we do not introduce any other potential so the lengthscales of the generated vortices are fixed by the scale of the cosmic string wake and the background magnetic field. Hence no turbulence occurs in our system.
@@
We plot the magnetic field lines in the simulation. The pressure density relation determines the epoch we are studying. Since this is the pre-recombination era, the magnetic field lines are deformed due to the flow of the fluid. We calculate the total magnetic field in the wake region as well as the peak value of the field. As long as the magnetic field remains constant, we do not expect any change in the magnetohydrodynamic flow. The flow remains steady with minor changes due to the velocity perturbation. We proceed to give a perturbation to the magnetic field. Since we would like the field to be amplified, we give a sheared perturbation to the magnetic field as it is previously known that sheared magnetic fields may lead to amplification of the magnetic field king . The field is perturbed first along the $x$- direction and then in the $y$-direction, the perturbations are denoted by $B_{x}$ and $B_{y}$, respectively. The results of these perturbations are discussed in the next section.
III Sheared magnetic fields and field amplification
@@
For a frozen in magnetic field, the frozen in condition ensures that the magnitude of the magnetic field will be related to the density of the plasma. It has been shown that an isotropic collapsing plasma can lead to the amplification of a magnetic field king . The argument was that the flux in the frozen in plasma is conserved, so in the collapsing region, as the cross-sectional area decreases, the magnitude of the magnetic field increases. Now, the formation of a wake behind a cosmic string leads to a substantial increase in the density of plasma behind the cosmic string. In this case too, there is the possibility that the magnitude of the magnetic field will be enhanced.
@@
To understand this, we looked at two different kinds of magnetic field perturbations. We start with an exponentially decreasing magnetic field perturbation. The strength of the magnetic field is proportional to the density of the magnetic field lines. As mentioned before, the cosmic string is moving in the $x$-direction. Due to the velocity perturbation, the fluid velocity has a component in the $y$-direction, and therefore, the magnetic field lines feel a force along the $y$-direction. Therefore, we start with a sheared perturbation in the direction of motion of the cosmic string. Sheared magnetic fields are usually generated by small scale density inhomogeneities and the lengthscale of the magnetic fields are therefore determined by the lengthscale of the inhomogeneities in the cosmic string wake. Fig.1 represents the graph of the total magnetic field with time for a perturbation of $B_{x}=B_{0}e^{-\alpha_{1}|y|}$. We call this a sheared magnetic field as the $x$ component of the magnetic field ($B_{x}$) at any given point has a variation in the $y$-direction. The lengthscale of the variation is given by the parameter $\alpha_{1}$. We have used $B_{0}=1$.
@@
Fig.2 represents the graph of the peak magnetic field evolution with time for $\alpha_{1}=0.01$. As the graphs shows there is an increase in the peak magnetic field value which will lead to the total magnetic energy being enhanced. The figure shows that the peak magnetic field value becomes at least five times the initial value. The peak magnetic value shown on the $y$- axis is the maximum value of the magnetic field at a specific point on the grid. Similarly, the total energy also becomes five times the initial energy value, the $|B|^{2}$ on the y - axis is the summation of the magnetic energy density over all the grid points. Together, they indicate that the magnetic field value not only increases at a particular point, it increases over the whole of the cosmic string wake. So, we see that the net magnetic field gets amplified.
The amplification is a result of the conical nature of the cosmic string metric. There is a velocity perturbation on the particles along the $y$-direction, so the particles move towards the conical space behind the cosmic string. We have given a sheared magnetic field as the magnetic field perturbation. Though we are perturbing the $x$ component, the variation is along the $y$-direction. As the magnetic field lines are being pushed in the $y$-direction by the velocity flow, so our perturbation is enhancing the effect of the compression of the magnetic field lines due to the fluid flow. We find that we get an amplification of the field for all values of $\alpha_{1}$.
This amplification can be attributed to the deformation tensor of the fluid element. The deformation tensor of the fluid element has been used to explain the amplification of frozen in magnetic fields in ref. king . The deformation parameter $D_{ij}$ represents the deformation of an infinitesimal fluid element of initial side length $\delta q_{j}$, which in this case is deformed due to the deficit angle in the local space of the cosmic string. The deformation tensor is defined by
@@
Fig.3 is used to explain how the magnetic field gets amplified for this given magnetic field perturbation. As has already been pointed out in ref. king , the magnitude of the magnetic field is affected only when the fluid motion is perpendicular to the magnetic field lines. In this case, for any given value of $x$, the magnetic field varies along $y$ as an exponentially decaying field. This is illustrated in Fig.3. The $q_{1}$ in the figure is the $x$-axis, while the $q_{2}$ is the $y$-axis. The angle $\theta$ by which the fluid element is deformed is proportional to the deficit angle of the cosmic string. The comoving volume element is preserved after the deformation, so, $\frac{V^{fin}}{V^{ini}}=D_{ij}$. Since the volume can be related to the density of the fluid, we will have
@@
where, $A$ denotes the initial and final cross-sectional areas through which the magnetic field line passes before and after deformation. If the length perpendicular to this area is referred to as $L$, from volume conservation we will have,
@@
The angle between $dq_{2}$ and $L^{fin}$ is the deficit angle $\theta$. This vector $L^{fin}$ has not been shown on the Fig.3 as it is rather difficult to depict in this case. Thus, we will have,
@@
This will mean that $B\propto\rho^{1/2}$. Similarly in any case, that is considered, the final magnetic field is determined by the final density of the deformed fluid. The deformation tensor we have discussed so far is also referred to as the Zel’dovich deformation tensor zeldovich and it depends on the Zel’dovich approximation. Therefore, since in the wake region, we should always have $\rho^{fin}>\rho^{ini}$, so the magnetic field will always be amplified in the wake of the cosmic string. This will happen as long as the magnetic field remains frozen in the plasma. We have checked the results for different values of $\alpha_{1}$ and in all the cases we see an increase in the magnitude of the magnetic field in the wake region.
The magnetic field in the cosmic string wake can have components in both the $x$-and $y$-direction. So, our next case would be to study a perturbation given by $B_{y}=B_{0}e^{-\alpha_{2}|(x-800)|}$. The origin is shifted to get a better figure of the final wake. It does not have anything to do with the field magnitude. Here, we depict the lengthscale by the parameter $\alpha_{2}$. The initial value is still $B_{0}=1$.
@@
Figure 5: Evolution of the peak value of the magnetic field in the cosmic string wake. The value saturates to about five times the initial value even in this case.
As is seen from Fig.4 and Fig.5, we still get a amplification of the magnetic field. The maximum value is approximately five times the initial value again. In this case too, the field lines are deformed by the perturbation. As we have taken the flow velocities and the perturbations along the $x$-and $y$-axes, the amplification is similar in both the cases. If the axes were not aligned, we might have got different results for the amplification. In both our cases the off-diagonal elements which are non-zero are driven by symmetry consideration. So, we can show that the magnetic field is proportional to the density of the wake in both the cases. We have repeated the simulation with a different periodic magnetic field and obtained similar results about the magnetic field amplification. For simplicity, we have taken one of the axes to be aligned with the flow, even if that is not the case, the amplification will still happen, only the symmetry consideration cannot be used to calculate the off-diagonal elements.
IV Magnetic field decay
@@
So, for a perturbation perpendicular to the flow velocity of the plasma, the lengthscale of the perturbation determines whether the field decays or it gets amplified. As discussed in the previous section, the Alfven’s theorem leads to the frozen in magnetic field lines. Due to the deformation of the fluid, the magnetic field lines get squeezed, and the amplification of the field is observed. However, in this case too the fluid is deformed but the magnetic field does not get amplified. This seems to indicate that Alfven’s theorem is no longer valid for this lengthscale. Zel’dovich approximation while true for all ideal fluids breaks down under certain conditions. The general assumption on which this is based is that the initial velocity field be irrotational. A magnetic field perturbation perpendicular to the flow will have a rotational component. If this rotational component is not negligible, the approximation will break down. Alfven’s theorem will also become invalid for such a flow.
@@
One of the ways to identify this lengthscale analytically, would be to study the particle trajectories of the charged particles in the plasma. For this, we start with obtaining the vector potential of the magnetic field perturbation. Though not unique, the most acceptable vector potential which will give us a magnetic field like $B_{y}=B_{0}e^{-\alpha_{2}|(x-800)|}$ is given by $\vec{A}=\hat{k}\frac{B_{0}}{\alpha_{2}}e^{-\alpha_{2}x}$. We have shifted the origin back to zero for ease of calculation. Since the vector component does not depend on the coordinates $(y,z)$, their respective momenta should be conserved. Therefore we can take the initial velocities as constant in these directions $v_{y}(t=0)=v_{0}$ and $v_{z}(t=0)=u_{0}$. Based on these constants, we can then obtain the effective potential which determines the trajectories of these particles. The effective potential is given by,
$\Phi(x)=\frac{\omega_{B}^{2}}{2\alpha_{2}^{2}}\left(e^{-2\alpha_{2}x}-1\right)+\frac{u_{0}\omega_{B}}{\alpha_{2}}\left(1-e^{-\alpha_{2}x}\right)$ (21)
Here $\omega_{B}$ is the gyrofrequency of the charged particle in the given magnetic field. This effective potential has an extremum point which will tell us whether the particle motion is bounded and stable or unbounded and unstable. If the particle motion is stable then the Alfven’s theorem would hold as the particle paths will be in well defined orbits and the mapping between the initial to the later coordinates remain single valued and continuous. For this the Zel’dovich’s approximation holds and the deformation tensor can be used to obtain the amplification of the field. When the particle path becomes unstable, Zel’dovich’s approximation no longer holds and the mapping may not be single valued. We will therefore obtain the condition for the extremum to be unstable to obtain the lengthscale at which the Alfven’s theorem breaks down. We find that the extremum is a minima given by,
$x_{min}=-\frac{1}{\alpha_{2}}log\left(-\frac{u_{0}\alpha_{2}}{\omega_{B}}\right)$ (22)
We assume that $\alpha_{2}$ being small, we can expand the logarithmic term in a series expansion. Considering only the first term of the series we find that $x_{min}\propto\frac{u_{0}}{\omega_{B}}$.
@@
So the minimum lengthscale for stability is proportional to the gyroradius as $\omega_{B}$ is the gyrofrequency of the charged particle. The gyroradius will be given by $r_{g}=\frac{mu_{0}}{qB}=\frac{u_{0}}{\omega_{B}}$. Assuming the majority of the particles to be electrons, one can then calculate the gyroradius of the plasma electrons. We find that the gyroradius gives us a good estimate of whether $l_{0}$ has been reached. Based on the fluid velocity, we calculate an approximate value of the electron gyroradius $r_{g}$. We find that the magnetic field gets amplified only when the particle gyroradius $r_{g}$ is smaller than the scale at which the magnetic field perturbation varies, i.e. for $r_{g}<1/\alpha_{2}$. Since in our numerical simulations the approximate value of the gyroradius is of the order of $r_{g}\sim 1.31$, hence we find that when the value of $\alpha_{2}\sim 1$, the field decays instead of getting amplified. This means that the particle motion is no longer stable and it does not have a well defined orbit anymore. Both Zel’dovich’s approximation as well as Alfven’s theorem break down as the mapping between the initial point in the path to the final point in the path of the particle no longer remains unique and single valued.
@@
Another explanation for this lengthscale can be obtained from the description of the motion of charged particles in terms of the gradient and curvature drift. This approach has been studied in specific sheared force-free magnetic fields vekstein before. In this approach, the motion of the electrically charged particles in a perturbed magnetic field has been studied using the approximate description called the guiding center or drift approximation northrop . The charged particle motion in the plasma is decomposed into a nearly circular fast gyromotion about a guiding center and the trajectory of the guiding center itself. Neither of the trajectories are constant for the non-uniform magnetic field. The overall motion is obtained by solving the second-order differential equations for the guiding center. Usually, for better understanding, the equations are separated into parallel and perpendicular components of the given magnetic field. In our present case, the perpendicular direction is the direction along which the field is non-uniform. The velocity in the perpendicular direction (i.e. along the non-uniform direction) is then given by a gradient drift ($v_{g}$) and a curvature drift ($v_{c}$). These are defined by,
@@
V Summary and Conclusions
In this work, we have studied the evolution of a magnetic field in a cosmic string wake. This is an important thing to understand as in recent times there is a renewed interest in identifying specific signatures of these strings through the generation of Gravitational Waves (GW)ligo . While it is possible that GW’s may lead to a signal for the cosmic string it is important to look for other signatures too. This is especially important as it has been recently established that cosmic string wakes may have magnetic fields within their wakes. Moreover since the width of the wake is quite narrow there is the possibility of magnetic reconnection. This means that apart from GW’s, synchrotron radiation as well as GRB’s may be additional features to look for when searching for these elusive objects. In view of this it is important to look at the lengthscales associated with the magnetic field evolution in the cosmic string wake.
@@ V Summary and Conclusions
In this work, we have found that the magnetic field in the wake region is amplified under certain circumstances of the flow. This is expected as the field lines are frozen in and so the conservation of magnetic flux density usually enhances the field when the fluid flow squeezes the flux lines together. The magnetic field is amplified due to the squeezing of the particle trajectories causing the overdensity in the wake region. Assuming the wake is symmetric about the flow axis, we show that the deformation tensor will have non-zero off-diagonal terms. This leads to the fact that the magnetic field strength will become proportional to the square root of the fluid density. Depending on the orientation of the deformation, the magnetic field can also be proportional to the density. Since the overdensity in the wake is always greater than the initial density of the flow, the magnetic field flux gets amplified in all the different cases. In this work, we use an exponentially decreasing magnetic field to illustrate our findings.
Our study also shows that if the sheared magnetic field perturbation in the cosmic string wake is aligned to the flow of the plasma past the string, then the magnetic field is always amplified for whatever perturbation is given. It is only when the magnetic field perturbation is perpendicular to the direction of the flow, the decay lengthscale of the perturbed magnetic field determines whether the field gets amplified or whether it decays. We study the lengthscales involved in these cases in detail. We find that the reason of the decay is the breakdown of the Zel’dovich approximation at lengthscale close to the gyroradius. We establish that this happens due to the fact that the effective potential generated by the moving electrons leads to an unstable orbit for the particle. Due to the unstable nature of the motion, the electron ceases to stay in a stable, bound orbit and instead moves away in random directions. In such a case, the Zel’dovich approximation fails as the initial and final points of the particles paths are no longer well defined.
@@
Thus we conclude that for a sheared perturbation which is perpendicular to the direction of flow of the fluid, the magnetic field lengthscale determines whether the magnetic field will amplify or decay. We have determined that the length scale below which the magnetic field decays is determined by the gyroradius of the charged particles being considered. If the lengthscale of the magnetic field is much larger than the gyroradius, Alfven’s theorem remains valid and we get the amplification of the field. However, if the lengthscale of the magnetic field perturbation is of the same order or less than the gyroradius of the electron then the field decays. We have presented results for a cosmic string moving in a high $\beta$ plasma as in the pre-recombination era the value of $\beta$ is high. The wake formation in our numerical study is also more pronounced for the high $\beta$ plasma. For the numerical part, we have used the OpenMHD code, an open source code, which is available from GitHub.
@@
The magnetic field due to the Biermann mechanism is due to the small-scale inhomogeneities generated in the cosmic string wake. Since the density perturbations of a long cosmic string are of varied lengthscales based on how they are generated, it is important to understand the behavior of the magnetic field for perturbations of different lengthscales. There are cosmic strings with small-scale structures on them which can generate small-scale non-linear density perturbations. It is also known that non-Gaussian density perturbations can be generated by long strings dacunha . As long as Alfven’s theorem holds, the magnetic field configurations and lengthscales will directly relate to these density perturbations. However, we have also found that for magnetic perturbations with lengthscales of the order of the gyro radius or less, Alfven’s theorem breaks down. This in turn can lead to the possibility of magnetic reconnection in the cosmic string wake. Since, in recent times, magnetic reconnection has been seen as a viable possibility in the narrow wakes of the cosmic string, this study will lead to a further understanding of the effect of these magnetic reconnection in the generation of GRB’s.
@@
Our study has some limitations as it is constrained to two dimensions only. The current study therefore has one dimensional sheared magnetic fields. The actual wake structure is three-dimensional, so it is also possible that the perturbation is a planar sheared field. The evolution of such perturbations may lead to further interesting results. We have also considered ideal MHD, it is quite possible that the reconnection possibility and the breakdown of Alfven’s theorem will be different for resistive MHD. In fact, there might be a higher probability of obtaining magnetic reconnection in that case. These and other aspects of magnetic field evolution in cosmic string wakes need to be studied in more details in the future.